Lab questions: I did some work but go over them because I’m not confident. Thanks.

Remember: report questions are individua! work Show your work for all calculations DATA AND CALCULATIONS Table A: Cylinder cylinder # set- Mass (g) Diameter (cm) Height (cm) Trial 1 Trial 2 Average 1,? 1-1 A1. Determine the denaity of the metal cylinder using the average data from Table A. The volume of a cylinder can be calculated as follows where h is the height, and d is the diameter, then calculate density: D mlv 221 2016766 Denaity: 7 bo g your calculated density, determine the volume (in in) of a cylinder weighing 29.2 ounces. Use dimensional for all steps. Table B: Irregular solid Unknown metal no.: Mass of solid (g) Volume of water (mL) Volume of water+ metal (mL) Volume of metal (mL) Trial 1 87 Trial 2 | 200Q 81. Determine the density of the metal using the appropriate data from Table B. Calculate the density for each trial, and then take the average. Be sure to include proper units. Trial 1: 1.303Trial 2 save this result for later in the semester, when we analyze these metals again (Expt 16) Trial 2: a GT3 165 Averaget 7.251
media%2Fdee%2Fdee266fc-d7e5-43c1-9477-20
Answer #1

A1) The volume of the cylinder is

V = ?h(d/2)2

where h = average height of the cylinder and d = average diameter of the cylinder. Plug in values and get

V = 3.14*(2.14 cm)*(1.27 cm/2)2 = 3.14*(2.14 cm)*(0.635 cm)2 = 2.70951 cm3.

The density of the cylinder is

d = M/V

where M = average mass of the cylinder. Plug in values and get

d = (20.76 g)/(2.70951 cm3) = 7.6619 g/cm3 ? 7.66 g/cm3 (ans).

A2) We know that

1 ounce = 28.3495 g.

1 inch = 2.54 cm.

Therefore, (1 inch)3 = 1 inch3 = (2.54 cm)3 = 16.38706 cm3.

The density of the cylinder is 7.66 g/cm3.

The volume of the cylinder is given as

V = M/d

Plug in values and obtain

V = (29.2 ounce)*(28.3495 g/1 ounce)/(7.66 gcm3) = 108.06895 cm3 = (108.06895 cm3)*(1 in3/16.38706 cm3) = 6.594773 in3 ? 6.59 in3 (ans).